OpenStudy (anonymous):

WILL MEDAL What is the solution to the equation 2(m-8)=4(m+6) ? m = ____

OpenStudy (jessereyna):

Hey i can solve it just give me a sec

OpenStudy (anonymous):

k

OpenStudy (f_jayyy):

2m-16 = 4m+24 -2m= 40 m= -20

OpenStudy (mathewr.):

m=-20

OpenStudy (annie12m):

2(M-8)=4(m+6) 2m-16=4m+24 6m-16=24 6m=30 m=5

OpenStudy (f_jayyy):

thats wrong because when you move the 4m to the other side you have to make it negative @annie12m

OpenStudy (annie12m):

oh ok sorry i did it wrong